OpenStudy (anonymous):

find the exact value of sin2π/3

OpenStudy (anonymous):

0.8660254038

OpenStudy (cwrw238):

sqrt3/2

OpenStudy (anonymous):

wat answer do u want

OpenStudy (anonymous):

find the coordinates of the point on the unit circle corresponding to the angle \(\frac{2\pi}{3}\) on the last page of the attached cheat sheet. the first coordinate is \(\cos(\frac{2\pi}{3}) \)the second coordinate is \(\sin(\frac{2\pi}{3})\)

OpenStudy (asnaseer):

\[\pi/3=60^0\]do you know the exact values of \(\sin(60)\) and \(\cos(60)\)? if you do then you can use the identity:\[\sin(2x)=2\sin(x)\cos(x)\]to work this out.

OpenStudy (anonymous):

too much work

OpenStudy (maheshmeghwal9):

k sir

OpenStudy (asnaseer):

you can also use the identity:\[\sin(x)=\sin(\pi-x)\]to get:\[\sin(2\pi/3)=\sin(\pi-2\pi/3)=\sin(\pi/3)\]

OpenStudy (maheshmeghwal9):

but why \[\sin(x)=\sin(\pi-x)\] It is becoz u should know that in 2nd quadrant sine is (+)ve.

OpenStudy (anonymous):

yes you can do either of those things, but it is really easier to locate \(\frac{2\pi}{3}\) on the unit circle and read off the coordinates

OpenStudy (maheshmeghwal9):

yes @satellite73 sir are correct even my words lie on the unit circle:)